OpenStudy (anonymous):

Find line perpendicular to 2x+3y=9 through the point (2,1)

OpenStudy (itsonlycdeee):

3y = -2x + 9 y = -2/3x + 3 perpendicular means the new line will be y = 3/2x + b. Now plug in the point (2,1) to find b & you have the new line.

OpenStudy (anonymous):

how do i plug in (2,1) @itsonlycdeee

OpenStudy (itsonlycdeee):

With (2, 1) . x = 2 & y = 1 1 = 3/2 (2) + b . Can you solve the rest from there?

OpenStudy (the_fizicx99):

Convert 2x+3y=9 into y = mx + b 2x+3y=9 Subtract both sides 2x 3y = -2x + 9 Divide every term by 3 to isolate y, y = -2/3x + 3 So now you have your point slope formula, now its asking for the perpendicular so its the reciprocal, -2/3 --> 3/2 y = 3/2x + 3 Your slope or M = 3/2. _____________________________ Now plug in the coordinates into: Y-y1 = M(x-x1) Y - 1=3/2(x-2) Distribute 3/2 to x and -2 y - 1 = 3/2x - 3 Add 1 to both sides, y = 3/2x -2 So that's your new equation

OpenStudy (anonymous):

Thank you both! @itsonlycdeee @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Np, you're welcome :3

OpenStudy (itsonlycdeee):

No problem!